In an engine cylinder a gas has a volumetric analysis of 13%CO2, 12.5%O2 and 74.5% N2. The temperature at the beginning of expansion is 9500C and gas mixture expands reversibly through a volume ratio of 8:1. According to the law pV1.2=constant. Find per kg of gas, the work done and the heat flow. Take Cp for CO2=1.235kJ/kg K and O2=1.088kJ/kg K and N2 is 1.172kJ/kg K.


1
Expert's answer
2021-01-26T13:16:03-0500

Percentage of mass (m) =

% of CO2=0.18

% of O2 = 0.13

% of N2 = 0.68

Specific heat at constant pressure "=0.18\times C_p1+0.13\times C_p2+0.68\times C_p3"

"C_p=1.17KJ\/Kg K"

Gas constant -

"\Rightarrow R=8.314(\frac{0.187}{0.44}+\frac{0.131}{0.44}+\frac{0.682}{0.44})"

"\Rightarrow R=0.2718 KJ/kg/K"

We know that,

"C_p-C_v=R"

"C_v=1.1725-0.271"

"C_v=0.9"

"i) W=\frac{R(T_1-T_2)}{n-1}"

"\frac{T_2}{T_1}=(\frac{v_1}{v_2})^{n-1}=(\frac{1}{8})^{1.2-1} =0.659"

W=566.8KJ/kg

"Q=\Delta U+W=U_2-U_1+W"

"=0.9\times 10^3(806-1223)+566.8\times 10^3"

"Q=191.41kJ\/kg"

"ii)" "S=R\log(\frac{v_2}{v_1}) =0.2718\times 10^3 \log 8"

"S=0.565KJ\/kg K"

"S=C_v \log(\frac{T_1}{T_2})"

"=0.9\times 10^3\log(\frac{122.3}{805.9})"

"=0.376"

Change in entropy "(S_2-S_1)=0.19kJ\/kg K"